# Nyström Kernel Stein Discrepancy

### v2

@@ A.2 Proof of Theorem 2
To conclude, we use decomposition (A.5), and union bound (A.7), (A.8), and (A.21). Further, we observe that $\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)=\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)$, and obtain that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|$ $\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}+$ (A.22)
$\quad+\sqrt{\frac{\lambda\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\log(12n/\delta)\log(12/\delta)}{m}}\Bigg{)}\geq 1-\delta$ (A.23)
provided that $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{\lambda},\log(3/\delta)\right\}$ and $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$ both hold. Now, specializing $\lambda=\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}$ for some absolute constant $c>0$, all constraints are satisfied for $m\gtrsim\max\left\{\log(3/\delta),\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}^{-1}\right\}$. Using our choice of $\lambda$, after rearranging, we get the stated claim.
@@ A.3 Proof of Corollary 1
  • Polynomial decay. The $\sqrt{n}$-consistency of the first two addends in Theorem 2 was established in the discussion following the statement. Hence, we limit our considerations to the last addend. Assume that $\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\lesssim\lambda^{-\gamma}$ for $\gamma\in(0,1]$. Observing that the trace expression is constant, the last addend in Theorem 2 yields that
    $\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}\right)}{m^{2}}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\sqrt{\frac{\log(12/\delta)\log(12n/\delta)}{m^{2-\gamma}}}\stackrel{{\scriptstyle(b)}}{{=}}\mathcal{O}\left(\frac{1}{\sqrt{n}}\right),$ (A.25)
@@
  • Exponential decay. Assume it holds that $\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\lesssim\log(1+c_{1}/\lambda)$. Observe that as per the discussion following Theorem 2, the first two addends are $\mathcal{O}\left(n^{-1/2}\right)$. For the last addend, again noticing that the trace is constant, we have
    $\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\mathcal{N}_{\mathbb{Q},h_{p}}\!\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}\right)}{m^{2}}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\log\left(1+\frac{c_{1}m}{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\right)}{m^{2}}}$ (A.26)
    $\stackrel{{\scriptstyle(b)}}{{\lesssim}}\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\right)}{m^{2}}}\stackrel{{\scriptstyle(c)}}{{=}}\mathcal{O}\left(\frac{1}{\sqrt{n}}\right),$ (A.27)
    where (a) uses the exponential decay assumption. (b) uses that $n\geq m$ and that the logarithm is a monotonically increasing function. (c) follows from our choice of $m\gtrsim\sqrt{n}\sqrt{\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\right)\log(12n/\delta)\log(12/\delta)}$, finishing the proof of the corollary.
@@ Appendix B Auxiliary results
This section collects our auxiliary results. Lemma B.1 builds on Rudi et al. [2015, Lemma 6], which assumes bounded feature maps, and on Della Vecchia et al. [2021, Lemma 5], which is stated in the context of leverage scores. Lemma B.2 states that a sub-exponential random variable satisfies Bernstein’s conditions, and Lemma B.3 is about the sub-Gaussianity of norms of Hilbert space-valued random variables. In Lemma B.4, we show how tensor products interplay with linearly transformed vectors. Lemma B.5 is about the maximum of real-valued sub-Gaussian random variables; it is a slightly altered restatement of Canonne [2021].
Lemma B.1.
Let Assumption 1 hold, and assume $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, for any $\delta\in(0,1)$, it holds that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\left(\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\lesssim\lambda\right)\geq 1-\delta$ (B.28)
provided that $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{\lambda},\log(1/\delta)\right\}$.
@@
Let the covariance of the Nyström samples be $\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}=\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\otimes h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)$, and $\beta(\lambda)=\lambda_{\text{max}}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},h_{p}}-\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right)$. The application of Rudi et al. [2015, Proposition 3 and Proposition 7] (recalled in Theorem D.6) yields the non-probabilistic inequality
@@
$\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)},$ (B.29)
when $\beta(\lambda)<1$. $\beta(\lambda)$ measures the deviation—in the sense of $\lambda_{\text{max}}(\cdot)$—of an empirical covariance to the population one. We will establish that our setting satisfies the conditions of Koltchinskii and Lounici [2017, Theorem 9] (recalled in Theorem D.3), with which we will show that the stronger requirement
$\beta(\lambda)\stackrel{{\scriptstyle(a)}}{{\leq}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},h_{p}}-\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}$ (B.30)
=$\stackrel{{\scriptstyle(b)}}{{=}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}-C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}$ (B.31)
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}\left\|{\frac{1}{m}\sum_{i=1}^{m}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\otimes h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)C_{\mathbb{Q},h_{p},\lambda}^{-1/2}-C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}<1\hskip 28.45274pt$ (B.32)
is satisfied for $m$ large enough. In (a), we used that the spectral radius is bounded by the operator norm. (b) is the distributive property, and (c) uses linearity.
In the following we prove (B.32). Let the Nyström selection be $\left(i_{j}\right)_{j=1}^{m}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Unif}([n])$; with this notation $\tilde{\mathbf{x}}_{j}=\mathbf{x}_{i_{j}}$ ($j\in[m])$. First, we condition on a fixed choice of $\left(i_{j}\right)_{j=1}^{m}$-s. Let us define the $Y=C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)\in\mathcal{H}_{h_{p}}$ random variable with $X\sim\mathbb{Q}$, and consider the samples $y_{i_{j}}:=C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{\mathbf{x}_{i_{j}}}\right)\in\mathcal{H}_{h_{p}}$. $Y$ has covariance
@@
$\mathbb{E}_{X\sim\mathbb{Q}}\left[Y\otimes Y\right]\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{X\sim\mathbb{Q}}\left[C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)\otimes h_{p}\left(\cdot,{X}\right)C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right]\stackrel{{\scriptstyle(b)}}{{=}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2},$ (B.33)
where (a) is implied by Lemma B.4 and (b) follows from Steinwart and Christmann [2008, (A.32)] by swapping the expectation with bounded linear operators. Similarly, we observe that the empirical covariance is $\frac{1}{m}\sum_{j=1}^{m}y_{i_{j}}\otimes y_{i_{j}}=C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$. In other words, this setup brings us into the setting of Theorem D.3, and it remains to show that $Y$ is sub-Gaussian in the sense of (2). Notice that $Y$ satisfies both conditions:
  • Mean-zero: By $\mathbb{P}=\mathbb{Q}$ and (6), $\mathbb{E}Y=C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\mathbb{E}h_{p}\left(\cdot,{X}\right)=0$.
  • (2) holds: Our goal now is to show that $\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ holds for all $u\in\mathcal{H}_{h_{p}}$. Let $u\in\mathcal{H}_{h_{p}}$ be arbitrary, and $v=C_{\mathbb{Q},h_{p},\lambda}^{-1/2}u\in\mathcal{H}_{h_{p}}$; $v$ is well-defined due to the invertibility of $C_{\mathbb{Q},h_{p},\lambda}$. Using this notation, we have
@@
    $\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(a)}}{{=}}\left\|{\left\langle{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(b)}}{{=}}\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),C_{\mathbb{Q},h_{p},\lambda}^{-1/2}u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}$ (B.34)
    =$\stackrel{{\scriptstyle(c)}}{{=}}\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),v}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(d)}}{{\lesssim}}\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),v}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(e)}}{{=}}\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),C_{\mathbb{Q},h_{p},\lambda}^{-1/2}u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}$ (B.35)
    =$\stackrel{{\scriptstyle(f)}}{{=}}\left\|{\left\langle{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(g)}}{{=}}\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(h)}}{{<}}\infty.$ (B.36)
    In (a), we use the definition of $Y$. (b) follows from the self-adjointness of $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$. (c) uses the definition of $v$. (d) is implied by Assumption 1(i) holding for all $v\in\mathcal{H}_{h_{p}}$. In (e) we again used the definition of $v$. (f) is implied by the self-adjointness of $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$. (g) is based on the definition of $Y$. (h) holds as $\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),v}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ by Assumption 1(i) and the expressions in-between were equalities. This establishes the sub-Gaussianity of $Y$.
Now, we can invoke Koltchinskii and Lounici [2017, Theorem 9] (recalled in Theorem D.3), and, by using (B.32), we obtain that with probability at least $1-\delta$ it holds that
$\beta(\lambda)\lesssim\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}\max\left\{\sqrt{\frac{r\left(C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right)}{m}},\sqrt{\frac{\log(1/\delta)}{m}}\right\}$ (B.37)
$=\max\Bigg{\{}\underbrace{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}^{1/2}}_{<1}\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)}{m}},$ (B.38)
$\qquad\qquad\underbrace{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}}_{<1}\sqrt{\frac{\log(1/\delta)}{m}}\Bigg{\}}$ (B.39)
provided that $m\geq\max\left\{r\left(C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right),\log(1/\delta)\right\}$, where $r(A):=\frac{\operatorname{tr}(A)}{\left\|{A}\right\|_{\mathrm{op}}}$.
We next simplify the requirements on $m$. By using that $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}=:\lambda_{1}$, we have that $\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}=\frac{\lambda_{1}}{\lambda_{1}+\lambda}\geq\frac{1}{2}$. Now also observing that
$\operatorname{tr}\left(C_{\mathbb{Q},h_{p}\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)\leq\frac{1}{\lambda}\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)$ (B.40)
and by upper bounding the nominator and lower bounding the denominator in $r$:

### v4

@@ A.2 Proof of Theorem 2
To conclude, we use decomposition (A.5), and union bound (A.7), (A.8), and (A.24). Further, we observe that $\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)=\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}(\lambda)$, and obtain that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|$ $\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)\log(6/\delta)}{n}}+$ (A.25)
$\quad+\sqrt{\frac{\lambda\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}(\lambda)\log(12n/\delta)\log(12/\delta)}{m}}\Bigg{)}\geq 1-\delta$ (A.26)
provided that $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},1\right\}\log(12/\delta)$ and $0<\lambda\leq\left\|{C_{\mathbb{Q},\bar{h}_{p}}}\right\|_{\mathrm{op}}$ both hold. Now, specializing $\lambda=\frac{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{m}$ for some absolute constant $c>1$, all constraints are satisfied for $m\gtrsim\max\left\{\log(12/\delta),\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)\left\|{C_{\mathbb{Q},\bar{h}_{p}}}\right\|_{\mathrm{op}}^{-1}\right\}$. Using our choice of $\lambda$, after rearranging, we get the stated claim.
@@ A.3 Proof of Corollary 1
  • Polynomial decay. The $\sqrt{n}$-consistency of the first two addends in Theorem 2 was established in the discussion following the statement. Hence, we limit our considerations to the last addend. Assume that $\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}(\lambda)\lesssim\lambda^{-\gamma}$ for $\gamma\in(0,1]$. Observing that the trace expression is constant, the last addend in Theorem 2 yields that
    $\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{m}\right)}{m^{2}}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\sqrt{\frac{\log(12/\delta)\log(12n/\delta)}{m^{2-\gamma}}}\stackrel{{\scriptstyle(b)}}{{=}}\mathcal{O}\left(\frac{1}{\sqrt{n}}\right),$ (A.28)
@@
  • Exponential decay. Assume it holds that $\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}(\lambda)\lesssim\log(1+c_{1}/\lambda)$. Observe that as per the discussion following Theorem 2, the first two addends are $\mathcal{O}\left(n^{-1/2}\right)$. For the last addend, again noticing that the trace is constant, we have
    $\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}\!\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{m}\right)}{m^{2}}}\stackrel{{\scriptstyle(a)}}{{\lesssim}}\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\log\left(1+\frac{c_{1}m}{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\right)}{m^{2}}}$ (A.29)
    $\stackrel{{\scriptstyle(b)}}{{\lesssim}}\sqrt{\frac{\log(12/\delta)\log(12n/\delta)\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\right)}{m^{2}}}\stackrel{{\scriptstyle(c)}}{{=}}\mathcal{O}\left(\frac{1}{\sqrt{n}}\right),$ (A.30)
    where (a) uses the exponential decay assumption. (b) uses that $n\geq m$ and that the logarithm is a monotonically increasing function. (c) follows from our choice of
    $m\gtrsim\sqrt{n}\sqrt{\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\right)\log(12n/\delta)\log(12/\delta)},$ (A.31)
    finishing the proof of the corollary.
A.4 Proof of Theorem 3
By the reverse triangle inequality, we obtain
$\left|S_{p}(\mathbb{Q})-S_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\leq\left\|{\mu_{h_{p}}(\mathbb{Q})-\mu_{h_{p}}\left(\hat{\mathbb{Q}}_{n}\right)}\right\|_{\mathcal{H}_{h_{\mathbb{P}}}}=\Bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\underbrace{\left[h_{p}(\cdot,X_{i})-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}(\cdot,X)\right]}_{=:\eta_{i}}\Bigg{\|}_{\mathcal{H}_{h_{p}}},$ (A.32)
which measures the concentration of i.i.d. centered random variables. To obtain the bound, we will use Bernstein’s inequality (recalled in Theorem C.4) by gaining control on the moments of $\left\|{\eta_{i}}\right\|_{\mathcal{H}_{hp}}$ with Lemma B.2.
First, note that the $\left\|{\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}$-s ($i\in[n]$) are sub-Gaussian as
$\left\|{\left\|{\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\left\|{\left\|{h_{p}\left(\cdot,{X_{i}}\right)-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}$ (A.33)
$\stackrel{{\scriptstyle(b)}}{{\leq}}\left\|{\left\|{h_{p}\left(\cdot,{X_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}+\left\|{\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}$ (A.34)
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left\|{\left\|{h_{p}\left(\cdot,{X_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}+\mathbb{E}_{X\sim\mathbb{Q}}\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(d)}}{{\lesssim}}\left\|{\left\|{h_{p}\left(\cdot,{X_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty.$ (A.35)
We use the definition of $\eta_{i}$ in (a). (b) is implied by the triangle inequality and the monotonicity of the norm. (c) is by Jensen’s inequality holding for Bochner integrals, and (d) follows from Lemma C.2(1); finiteness is due to the imposed assumption.
Hence, $\left\|{\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}$ is sub-exponential (Lemma C.2(3)), and, by Lemma B.2, it holds for any $p\geq 2$ that
$\mathbb{E}_{X\sim\mathbb{Q}}\left\|{\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}^{p}\leq\frac{1}{2}p!\sigma^{2}B^{p-2},$ (A.36)
with $\sigma,B\lesssim\left\|{\left\|{\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{1}}=:K$. Now, applying Theorem C.4 yields that, for any $\delta\in(0,1)$, it holds with probability at least $1-\delta$ that
$\left\|{\frac{1}{n}\sum_{i=1}^{n}\eta_{i}}\right\|_{\mathcal{H}_{h_{p}}}\lesssim\frac{2K\log(2/\delta)}{n}+\sqrt{\frac{2K^{2}\log(2/\delta)}{n}},$ (A.37)
which is the stated claim.
@@ Appendix B AUXILIARY RESULTS
This section collects our auxiliary results. Lemma B.1 builds on Rudi et al. [2015, Lemma 6], which assumes bounded feature maps, and on Della Vecchia et al. [2021, Lemma 5], which is stated in the context of leverage scores. The main technical challenge that we resolve lies in introducing and handling the centered covariance operator that allows us to make use of existing concentration results. Lemma B.2 states that a sub-exponential random variable satisfies Bernstein’s conditions, and Lemma B.3 is about the sub-Gaussianity of norms of Hilbert space-valued random variables. In Lemma B.4, we show how tensor products interplay with linearly transformed vectors. Lemma B.5 is about the maximum of real-valued sub-Gaussian random variables; it is a slightly altered restatement of Canonne [2021]. In Lemma B.6 and Lemma B.7, we collect inequalities of positive operators and of norms of covariance operators, respectively.
Lemma B.1 (Projected covariance operator bound).
Let Assumption 1 hold, and assume $0<\lambda\leq\left\|{C_{\mathbb{Q},\bar{h}_{p}}}\right\|_{\mathrm{op}}$. Then, for any $\delta\in(0,1)$, it holds that
$\left(\mathbb{P}^{n}\otimes\Lambda^{m}\right)\left(\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\lesssim\lambda\right)\geq 1-\delta,$ (B.38)
provided that $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},1\right\}\log\left(4/\delta\right)$.
@@
The proof proceeds in two steps: First, we show that $\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)}$, when $\beta(\lambda):=\lambda_{\text{max}}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\right)<1$, where
@@
$\tilde{h}_{p}(\cdot,\mathbf{x})$ $:=h_{p}\left(\cdot,{\mathbf{x}}\right)-\frac{1}{m}\sum_{i\in[m]}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\quad(\mathbf{x}\in\mathbb{R}^{d}),$ (B.39)
@@
$C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p}}$ $=\frac{1}{m}\sum_{i\in[m]}\tilde{h}_{p}\left(\cdot,\tilde{\mathbf{x}}_{i}\right)\otimes\tilde{h}_{p}\left(\cdot,\tilde{\mathbf{x}}_{i}\right)$ (B.40)
$=\frac{1}{m}\sum_{i\in[m]}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\otimes h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)-\left(\frac{1}{m}\sum_{i\in[m]}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\right)\otimes\left(\frac{1}{m}\sum_{i\in[m]}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\right).$ (B.41)
In the second step, we show that $\beta(\lambda)<1$ (with high probability) for $m$ large enough.
Step 1.
@@
$\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\lambda\left\|{\left(Z^{*}_{m}Z_{m}+\lambda I\right)^{-1/2}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(b)}}{{=}}\lambda\left\|{C_{\tilde{\mathbb{Q}}_{m},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}$ (B.42)